OpenStudy (anonymous):

four-fifths of what number is 7/2?

OpenStudy (freckles):

set up the equation ... \[\frac{4}{5}n=\frac{7}{2} \\ \text{ or } \frac{4n}{5}=\frac{7}{2}\]

OpenStudy (freckles):

can you solve this ?

OpenStudy (anonymous):

I can try

OpenStudy (anonymous):

would I multiply both sides